### v3

@@ 3.3 Volume Rendering and Super-resolution
The tri-plane can be rendered into a 2D RGB image through the differentiable volume rendering [40, 36, 10]. The expression-transferred tri-plane $\text{T}_{\beta_{D}}(S)$ is projected onto 3 orthogonal planes ($xy,yz,zx$-planes) and then aggregated through average [10]:
@@
The differentiable volume rendering [40, 10, 22] composites each color $c$ and density $\mathbf{\sigma}$ into a RGB value $\mathcal{C}$ along the camera ray $\mathbf{r}$:
[…]
@@
Directly rendering a target high-resolution image requires high computational cost. One promising approach to address this issue is to incorporate super-resolution blocks [10, 61, 56] that upsamples the rendered image of low resolution. Following this approach, we first render a $\hat{D}_{raw}\in\mathbb{R}^{3\times\frac{H}{4}\times\frac{W}{4}}$ and then apply the super-resolution to obtain the target resolution $\hat{D}\in\mathbb{R}^{3\times H\times W}$, as illustrated in Fig. 1. Instead of using style-modulated convolution [10, 56], we use plane convolutional blocks for super-resolution, as we do not leverage the style latent code. Detailed architecture is provided in supplementary materials.
@@ 4 Experiments
Table 1: Quantitative comparison on VFHQ. The best score for each metric is in bold. Note that we only measure CSIM [14], AED and APD [16, 47] for the cross-identity experiment as no ground-truth is available.
@@
StyleHEAT [65] 14.233 0.428 30.406 0.464 0.161 0.139 0.505 0.242 0.136
DPE [42] 23.241 0.750 3.661 0.831 0.083 0.032 0.586 0.253 0.085
ROME [30] 14.185 0.642 7.281 0.737 0.111 0.051 0.641 0.224 0.074
OTAvatar [38] 17.441 0.651 11.502 0.662 0.176 0.067 0.610 0.290 0.198
HiDe-NeRF [34] 21.228 0.728 8.245 0.867 0.106 0.049 0.707 0.255 0.065
@@ 4.1 Dataset and Pre-processing
We train our model on real video dataset VFHQ [64]. Following the video pre-processing strategies in [52, 31], we convert the original video into 25 fps and crop the facial regions of resolution $256\times 256$, ensuring that the nose is located at the center of the image. We use a 3DMM extractor [16] to obtain the expression parameter $\beta\in\mathbb{R}^{64}$. We adopt the pre-preprocesing strategy of EG3D [10] for the camera parameter $p\in\mathbb{R}^{25}$ (the concatenation of the camera intrinsic parameters in $\mathbb{R}^{9}$ and the inverse extrinsic parameter in $\mathbb{R}^{16}$). After the video pre-processing, 6196 video clips are used for training, and 50 videos are used for test. We also evaluate our model on the test dataset of TalkingHead-1KH [58]. After the same pre-processing, remaining 20 videos of different identities are used.
@@ 4.2 Evaluation
Table 2: Quantitative comparison on TalkingHead-1KH. The best score for each metric is in bold. Note that we only measure CSIM [14], AED and APD [16, 47] for the cross-identity experiment as no ground-truth is available.
@@
StyleHEAT [65] 15.613 0.517 21.198 0.575 0.148 0.095 0.571 0.218 0.102
DPE [42] 23.201 0.786 4.281 0.807 0.093 0.029 0.714 0.216 0.081
ROME [30] 15.921 0.695 13.444 0.726 0.123 0.062 0.667 0.201 0.084
OTAvatar [38] 16.952 0.660 11.615 0.668 0.181 0.063 0.682 0.247 0.150
HiDe-NeRF [34] 19.759 0.729 5.746 0.843 0.112 0.043 0.757 0.232 0.085
@@
We compare our model against 2D-based [65, 42] and 3D-based [30, 38, 34] image-driven portrait animation methods whose official implementations are available. StyleHEAT [65] warps the 2D spatial features of pre-trained StyleGAN2 using 3DMM parameters, DPE [42] disentangles the pose and the expression in the motion latent space without using 3DMM parameters. ROME [30] is a mesh-based method transferring the expression and pose using 3DMM. OTAvatar [38] leverages pre-trained EG3D [10] by modeling head motion in terms of latent codes. HiDe-NeRF [34] deforms the source tri-plane by predicting expression-aware deformation fields. For evaluation, we employ peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) for image quality, average key-point distance (AKD) [47] for facial structure based on the 68 facial key-points, cosine similarity of identity embedding (CSIM) [14] for identity preservation, average expression distance (AED), and average pose distance (APD) [16, 47] for expression transferring and pose matching. For the cross-identity experiments, we only measure CSIM, AED and APD as no ground-truth image is available.
Table 3: Ablation studies on the expression encoding. Same evaluation setting with Tab. 1. The best score for each metric is in bold.
@@
Same-identity experiments.  We report the same-identity transfer experiment results in Tab. 1 and Tab. 2, and illustrate the qualitative results in Fig. 4. For a fair comparison, ROME [30], OTAvatar [38], and HiDe-NeRF [38] are evaluated on the foreground facial region with different field of view. DPE [42] shows the stable performance in the same-identity experiments with the fine-grained expression controls. Among the 3D-based methods, HiDe-NeRF [34] scores the highest in the identity preservation (CSIM). Our method scores the best result in the majority of evaluation metrics. Especially, it has an advantage in expression controls (AKD and AED).
Cross-identity experiments.  In Tab. 1 and Tab. 2, we also conduct the cross-identity transfer experiments that transfers the expression and pose of different identity into the source identity. As illustrated in Fig. 5, DPE [42] shows visual artifacts and appearance swap, such as face contours and eye shape, due to the insufficient disentanglement of expression and pose in the motion space. HiDe-NeRF [34] scores the highest identity preservation (CSIM) while un-predictable light changes are involved due to the point-wise deformation field on the tri-plane. Our method can transfer the driving expression without appearance swap and generates a video without video-level artifacts such as light changes and flickers. Please refer to our supplementary videos.
@@ 4.3 Ablation Studies and Further Results
Ablation studies on the expression encoding.  In Tab. 3, we conduct ablation studies on different expression encoding strategies. In Direct 3DMM, we replace our CLeBS with fully-connected layers to directly inject the expression parameters of 3DMM through EAdaLN. As illustrated in Fig. 6, the direct injection does not change appearance when transferring same-identity expression however, it changes appearance (e.g., eyebrows and facial contour) when transferring cross-identity expression. Furthermore, since the raw expression parameters inherently contain noise, the generated image also exhibits visual artifacts. In E2E LeBS, we decrease the the number of basis vectors $n$ in LeBS for appearance bottleneck to validate the proposed contrastive pre-training. Each LeBS with $n=25,10,5$ is jointly trained (i.e., E2E) with entire model without any pre-training. Due to the entanglement of appearance and expression, both appearance and expression are changed as a whole as the the number of basis vector $n$ decreases. LeBS alone is insufficient for extracting appearance-free expression from the expression parameters.
Ablation studies on EAdaLN.  In Tab. 4, we conduct ablation studies on EAdaLN (w. EAdaLN) by comparing it with cross-attention (w. Cross-attention), which is a widely used conditioning method in transformer-based architectures [57, 44]. Specifically, we replace all the EAdaLN blocks in $\mathbf{G}$ (Fig. 3) with cross-attention blocks. In both scenarios, CLeBS serves the refined expression $\beta^{\prime}$. As shown in Tab. 4 and Fig. 7, the cross-attention fails to handle the expression accurately, which verifies the effectiveness of our EAdaLN for the expression conditioning.
Visualization of facial expression parameters.  In Fig. 10, we sample 10 random frames from 10 different videos of distinct individual in VFHQ [64] and visualize the low-dimensional t-SNE [39] results of the two expression parameters: $\beta\in\mathbb{R}^{64}$ and $\beta^{\prime}\in\mathbb{R}^{d}$. In Fig. 8(a), the 3DMM expression parameters show strong entanglement with respective to their identities, indicating the hidden appearance information in them. On the other hand, as shown in Fig. 8(b), our contrastive pre-training mitigates the entanglement, thereby resolving the appearance swap in the cross-identity expression transfer in Fig. 6.
@@
Figure 8: Visualization of the expression parameters. We plot t-SNE [39] of raw 3DMM expression and our appearance-free expression parameter.
@@
Figure 10: Novel-view synthesis results with expression transfer. Our method can generate more multi-view consistent images compaired to HiDe-NeRF [34].
Linear scaling along the orthogonal directions.  In Fig. 10, we verify that $\beta^{\prime}\in\mathbb{R}^{d}$ has the orthogonal structure where the learned basis $\mathbf{V}$ determines the different expressions even if trained in unsupervised manner and the coefficients $\{\lambda_{i}\}_{i=1}^{n}$ scale their intensities. Specifically, we visualize two orthogonal directions $\mathbf{v}_{4}$ and $\mathbf{v}_{8}$ and linearly scale their coefficients $\lambda_{4}$ and $\lambda_{8}$ from $1$ to $10$. As shown in Fig. 8(c), $\mathbf{v}_{4}$ controls eye closing and mouth closing, while Fig. 8(d) illustrates that $\mathbf{v}_{8}$ controls mouth opening. Notably, the orthogonal basis does not influence head movements.
Novel-view synthesis with expression transfer.  In Fig. 10, we compare the results of novel-view synthesis with expression transfer to those of HiDe-NeRF [34]. Both methods utilize the tri-plane and differentiable volume rendering to generate novel-view images. However, while HiDe-NeRF transfers the driving expression by deforming the generated tri-plane into a canonical tri-plane based on driving conditions [43], our method relies on the hybrid generator $\mathbf{G}$ with EAdaLN. In both same-identity and cross-identity transfer scenarios, our method synthesizes more view-consistent results, highlighting the effectiveness of our method in expression transfer without relying on deformation. Please refer to supplementary videos.
@@ 5 Conclusion
Limitations and future work.  While our method can generate realistic face videos with driving expressions and views, it still has several limitations. First, our method cannot separately generate the foreground and background regions as the tri-plane representation construct them as a whole. Several works address this limitation by extending the tri-plane representation [4], restricting rendering points in the ray marching process [51], or leveraging the off-the-shelf segmentation model [29] to manually separate them [34, 35, 38]. Second, our method cannot control non-facial body parts (e.g., neck and shoulders) and eye gazing as they are beyond the capability of the 3DMM parameters. We plan to address these limitations for future work.

### v4

@@ 3.3 Volume Rendering and Super-resolution
The tri-plane can be rendered into a 2D RGB image through the differentiable volume rendering [nerf, neuralvolumes, eg3d]. The expression-transferred tri-plane $\text{T}_{\beta_{D}}(S)$ is projected onto 3 orthogonal planes ($xy,yz,zx$-planes) and then aggregated through average [eg3d]:
@@
The differentiable volume rendering [nerf, eg3d, adnerf] composites each color $c$ and density $\mathbf{\sigma}$ into a RGB value $\mathcal{C}$ along the camera ray $\mathbf{r}$:
$\mathcal{C}=\int_{t_{n}}^{t_{f}}\mathbf{\sigma}(\mathbf{r}(t))\cdot c(\mathbf{r}(t))\cdot T(t)dt,$ (9)
@@
Directly rendering a target high-resolution image requires high computational cost. One promising approach to address this issue is to incorporate super-resolution blocks [eg3d, gramhd, live3d] that upsamples the rendered image of low resolution. Following this approach, we first render a $\hat{D}_{raw}\in\mathbb{R}^{3\times\frac{H}{4}\times\frac{W}{4}}$ and then apply the super-resolution to obtain the target resolution $\hat{D}\in\mathbb{R}^{3\times H\times W}$, as illustrated in Fig.˜1. Instead of using style-modulated convolution [eg3d, live3d], we use plane convolutional blocks for super-resolution, as we do not leverage the style latent code. Detailed architecture is provided in supplementary materials.
@@ 4 Experiments
Table 1: Quantitative comparison on VFHQ. The best score for each metric is in bold. Note that we only measure CSIM [arcface], AED and APD [bfm, pirenderer] for the cross-identity experiment as no ground-truth is available.
@@
StyleHEAT [styleheat] 14.233 0.428 30.406 0.464 0.161 0.139 0.505 0.242 0.136
DPE [dpe] 23.241 0.750 3.661 0.831 0.083 0.032 0.586 0.253 0.085
ROME [rome] 14.185 0.642 7.281 0.737 0.111 0.051 0.641 0.224 0.074
OTAvatar [otavatar] 17.441 0.651 11.502 0.662 0.176 0.067 0.610 0.290 0.198
HiDe-NeRF [hidenerf] 21.228 0.728 8.245 0.867 0.106 0.049 0.707 0.255 0.065
@@ 4.1 Dataset and Pre-processing
We train our model on real video dataset VFHQ [vfhq]. Following the video pre-processing strategies in [fomm, stylelipsync], we convert the original video into 25 fps and crop the facial regions of resolution $256\times 256$, ensuring that the nose is located at the center of the image. We use a 3DMM extractor [bfm] to obtain the expression parameter $\beta\in\mathbb{R}^{64}$. We adopt the pre-preprocesing strategy of EG3D [eg3d] for the camera parameter $p\in\mathbb{R}^{25}$ (the concatenation of the camera intrinsic parameters in $\mathbb{R}^{9}$ and the inverse extrinsic parameter in $\mathbb{R}^{16}$). After the video pre-processing, 6196 video clips are used for training, and 50 videos are used for test. We also evaluate our model on the test dataset of TalkingHead-1KH [osfv]. After the same pre-processing, remaining 20 videos of different identities are used.
@@ 4.2 Evaluation
Table 2: Quantitative comparison on TalkingHead-1KH. The best score for each metric is in bold. Note that we only measure CSIM [arcface], AED and APD [bfm, pirenderer] for the cross-identity experiment as no ground-truth is available.
@@
StyleHEAT [styleheat] 15.613 0.517 21.198 0.575 0.148 0.095 0.571 0.218 0.102
DPE [dpe] 23.201 0.786 4.281 0.807 0.093 0.029 0.714 0.216 0.081
ROME [rome] 15.921 0.695 13.444 0.726 0.123 0.062 0.667 0.201 0.084
OTAvatar [otavatar] 16.952 0.660 11.615 0.668 0.181 0.063 0.682 0.247 0.150
HiDe-NeRF [hidenerf] 19.759 0.729 5.746 0.843 0.112 0.043 0.757 0.232 0.085
@@
We compare our model against 2D-based [styleheat, dpe] and 3D-based [rome, otavatar, hidenerf] image-driven portrait animation methods whose official implementations are available. StyleHEAT [styleheat] warps the 2D spatial features of pre-trained StyleGAN2 using 3DMM parameters, DPE [dpe] disentangles the pose and the expression in the motion latent space without using 3DMM parameters. ROME [rome] is a mesh-based method transferring the expression and pose using 3DMM. OTAvatar [otavatar] leverages pre-trained EG3D [eg3d] by modeling head motion in terms of latent codes. HiDe-NeRF [hidenerf] deforms the source tri-plane by predicting expression-aware deformation fields. For evaluation, we employ peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) for image quality, average key-point distance (AKD) [pirenderer] for facial structure based on the 68 facial key-points, cosine similarity of identity embedding (CSIM) [arcface] for identity preservation, average expression distance (AED), and average pose distance (APD) [bfm, pirenderer] for expression transferring and pose matching. For the cross-identity experiments, we only measure CSIM, AED and APD as no ground-truth image is available.
Table 3: Ablation studies on the expression encoding. Same evaluation setting with Tab.˜1. The best score for each metric is in bold.
@@
Same-identity experiments.  We report the same-identity transfer experiment results in Tab.˜1 and Tab.˜2, and illustrate the qualitative results in Fig.˜4. For a fair comparison, ROME [rome], OTAvatar [otavatar], and HiDe-NeRF [otavatar] are evaluated on the foreground facial region with different field of view. DPE [dpe] shows the stable performance in the same-identity experiments with the fine-grained expression controls. Among the 3D-based methods, HiDe-NeRF [hidenerf] scores the highest in the identity preservation (CSIM). Our method scores the best result in the majority of evaluation metrics. Especially, it has an advantage in expression controls (AKD and AED).
Cross-identity experiments.  In Tab.˜1 and Tab.˜2, we also conduct the cross-identity transfer experiments that transfers the expression and pose of different identity into the source identity. As illustrated in Fig.˜5, DPE [dpe] shows visual artifacts and appearance swap, such as face contours and eye shape, due to the insufficient disentanglement of expression and pose in the motion space. HiDe-NeRF [hidenerf] scores the highest identity preservation (CSIM) while un-predictable light changes are involved due to the point-wise deformation field on the tri-plane. Our method can transfer the driving expression without appearance swap and generates a video without video-level artifacts such as light changes and flickers. Please refer to our supplementary videos.
@@ 4.3 Ablation Studies and Further Results
Ablation studies on the expression encoding.  In Tab.˜3, we conduct ablation studies on different expression encoding strategies. In Direct 3DMM, we replace our CLeBS with fully-connected layers to directly inject the expression parameters of 3DMM through EAdaLN. As illustrated in Fig.˜6, the direct injection does not change appearance when transferring same-identity expression however, it changes appearance (e.g., eyebrows and facial contour) when transferring cross-identity expression. Furthermore, since the raw expression parameters inherently contain noise, the generated image also exhibits visual artifacts. In E2E LeBS, we decrease the the number of basis vectors $n$ in LeBS for appearance bottleneck to validate the proposed contrastive pre-training. Each LeBS with $n=25,10,5$ is jointly trained (i.e., E2E) with entire model without any pre-training. Due to the entanglement of appearance and expression, both appearance and expression are changed as a whole as the the number of basis vector $n$ decreases. LeBS alone is insufficient for extracting appearance-free expression from the expression parameters.
Ablation studies on EAdaLN.  In Tab.˜4, we conduct ablation studies on EAdaLN (w. EAdaLN) by comparing it with cross-attention (w. Cross-attention), which is a widely used conditioning method in transformer-based architectures [transformer, dit]. Specifically, we replace all the EAdaLN blocks in $\mathbf{G}$ (Fig.˜3) with cross-attention blocks. In both scenarios, CLeBS serves the refined expression $\beta^{\prime}$. As shown in Tab.˜4 and Fig.˜7, the cross-attention fails to handle the expression accurately, which verifies the effectiveness of our EAdaLN for the expression conditioning.
Visualization of facial expression parameters.  In Fig.˜10, we sample 10 random frames from 10 different videos of distinct individual in VFHQ [vfhq] and visualize the low-dimensional t-SNE [tsne] results of the two expression parameters: $\beta\in\mathbb{R}^{64}$ and $\beta^{\prime}\in\mathbb{R}^{d}$. In Fig.˜8(a), the 3DMM expression parameters show strong entanglement with respective to their identities, indicating the hidden appearance information in them. On the other hand, as shown in Fig.˜8(b), our contrastive pre-training mitigates the entanglement, thereby resolving the appearance swap in the cross-identity expression transfer in Fig.˜6.
@@
Figure 8: Visualization of the expression parameters. We plot t-SNE [tsne] of raw 3DMM expression and our appearance-free expression parameter.
@@
Figure 10: Novel-view synthesis results with expression transfer. Our method can generate more multi-view consistent images compaired to HiDe-NeRF [hidenerf].
Linear scaling along the orthogonal directions.  In Fig.˜10, we verify that $\beta^{\prime}\in\mathbb{R}^{d}$ has the orthogonal structure where the learned basis $\mathbf{V}$ determines the different expressions even if trained in unsupervised manner and the coefficients $\{\lambda_{i}\}_{i=1}^{n}$ scale their intensities. Specifically, we visualize two orthogonal directions $\mathbf{v}_{4}$ and $\mathbf{v}_{8}$ and linearly scale their coefficients $\lambda_{4}$ and $\lambda_{8}$ from $1$ to $10$. As shown in Fig.˜8(c), $\mathbf{v}_{4}$ controls eye closing and mouth closing, while Fig.˜8(d) illustrates that $\mathbf{v}_{8}$ controls mouth opening. Notably, the orthogonal basis does not influence head movements.
Novel-view synthesis with expression transfer.  In Fig.˜10, we compare the results of novel-view synthesis with expression transfer to those of HiDe-NeRF [hidenerf]. Both methods utilize the tri-plane and differentiable volume rendering to generate novel-view images. However, while HiDe-NeRF transfers the driving expression by deforming the generated tri-plane into a canonical tri-plane based on driving conditions [nerfies], our method relies on the hybrid generator $\mathbf{G}$ with EAdaLN. In both same-identity and cross-identity transfer scenarios, our method synthesizes more view-consistent results, highlighting the effectiveness of our method in expression transfer without relying on deformation. Please refer to supplementary videos.
@@ 5 Conclusion
Limitations and future work.  While our method can generate realistic face videos with driving expressions and views, it still has several limitations. First, our method cannot separately generate the foreground and background regions as the tri-plane representation construct them as a whole. Several works address this limitation by extending the tri-plane representation [panohead], restricting rendering points in the ray marching process [ballgan], or leveraging the off-the-shelf segmentation model [modnet] to manually separate them [hidenerf, goavatar, otavatar]. Second, our method cannot control non-facial body parts (e.g., neck and shoulders) and eye gazing as they are beyond the capability of the 3DMM parameters. We plan to address these limitations for future work.